OpenStudy (anonymous):

Which property of equality would be used to isolate the variable a in this equation? a-4=12

OpenStudy (anonymous):

if we perform the same function on the sides of an equality it will persist that is if we have x=y, we can always say that f(x)=f(y), so if we perform the function f(x)=x+4 on both sides we get => f(a-4)=f(12) a-4+4=12+4 a=16